# Collaborative Multi-Agent Optimization for Personalized Memory System

### 13 Mar 2026 v1

@@ Abstract
Memory systems are crucial to personalized LLMs by mitigating the context window limitation in capturing long-term user-LLM conversations. Typically, such systems leverage multiple agents to handle multi-granular memory construction and personalized memory retrieval tasks. To optimize the system, existing methods focus on specializing agents on their local tasks independently via prompt engineering or fine-tuning. However, they overlook cross-agent collaboration, where independent optimization on local agents hardly guarantees the global system performance. To address this issue, we propose a Collaborative Reinforcement Learning Framework for Multi-Agent Memory Systems (CoMAM), jointly optimizing local agents to facilitate collaboration. Specifically, we regularize agents’ execution as a sequential Markov decision process (MDP) to embed inter-agent dependencies into the state transition, yielding both local task rewards (e.g., information coverage for memory construction) and global rewards (i.e., query-answer accuracy). Then, we quantify each agent’s contribution via group-level ranking consistency between local and global rewards, treating them as adaptive weights to assign global credit and integrate local-global rewards. Each agent is optimized by these integrated rewards, aligning local improvements with the global performance. Experiments show CoMAM outperforms leading memory systems, validating the efficacy of our proposed collaborative reinforcement learning for joint optimization.
Multi-agent, Memory System, Reinforcement Learning
@@ 1 Introduction
Figure 1: Construction and retrieval agents, optimized on local tasks independently, yield lower global system performance than those under joint optimization.
Memory systems (Chhikara et al., 2025; Zhong et al., 2024; Hu et al., 2025) have emerged as a pivotal frontier for personalized large language models (LLMs), which preserve users’ long-term conversational history beyond LLMs’ context window and selectively integrate it to tailor personalized responses (Chen et al., 2024a; Zhang et al., 2025). Typically, the memory system comprises multiple specialized agents: several construction agents summarize and store user history into multi-granular memories (e.g., long/short or coarse/fine), respectively (Li et al., 2025; Wang & Chen, 2025; Wang et al., 2025); and retrieval agents select and integrate relevant memories for subsequent queries via RAG (Lewis et al., 2020) or search-based reasoning (Yu et al., 2026; Jin et al., 2025). To enhance system effectiveness, existing methods focus on optimizing these agents for their respective tasks independently: Prompt engineering approaches (Li et al., 2025; Zhong et al., 2024) specialize each agent for its local task with well-crafted prompts, while reinforcement learning methods (Wang et al., 2025; Zhou et al., 2026; Yan et al., 2025b) train agents separately using rewards such as the query-answer accuracy.
Despite local successes, they overlook cross-agent collaboration, treating the memory system as a “collection of local experts” rather than a collaborative team that evolves together for a global goal. Critically, independent optimization on local agents hardly guarantees global system performance (Akella, 2025; Wu et al., 2026). While individual agents may excel at their respective tasks, their independent strategies can conflict when combined in the global pipeline. To be more specific, if the fine-grained construction agent is optimized by treating retrieval agents as fixed environment components, it may store redundant or noisy content that impairs retrieval effectiveness. Conversely, if retrieval agents are trained without considering construction strategies, they may over-filter information and miss critical details during retrieval. As illustrated in Figure 1, even with independently well-trained agents (rewards converged), the combined system consistently underperforms that with joint optimization across different conversational history lengths. This motivates us to explore jointly optimizing multiple agents in the memory system to facilitate collaboration, which faces two challenges (see Figure 2):
Figure 2: Illustration of challenges for joint optimization.
  • Heterogeneity and Asynchrony: Multi-granular construction and retrieval agents operate with heterogeneous configurations and asynchronous executions across different tasks. Their discrete parameter space and inherent structural complexity complicate end-to-end optimization.
  • Local-Global Alignment: Fostering collaboration requires aligning local improvement with the global performance. However, directly assigning equal shares of the global objective to individual agents introduces ambiguity in task execution, as agents’ contributions are divergent.
@@
To address these challenges, we introduce CoMAM, a Collaborative reinforcement learning framework for Multi-Agent Memory systems. Specifically, we formulate a standard memory system with three specialized agents for multi-granular memory construction and retrieval: an extraction agent to construct fine-grained memories (e.g., key events and facts), a profile agent to construct coarse-grained memories (e.g., users’ personalized preferences), and a retrieval agent to retrieve memories and answer queries. To optimize these heterogeneous agents jointly, we frame their asynchronous execution as a sequential Markov decision process (MDP), which embeds inter-agent dependencies into state transitions and yields both local task-specific rewards (e.g., information coverage for memory construction) and global system rewards (i.e., query answer accuracy). To achieve local-global alignment, CoMAM quantifies each agent’s system contribution by measuring group-level ranking consistency between its local and global rewards, then adaptively allocates global rewards to agents based on this consistency. Each agent is ultimately optimized with a combination of its task-specific reward and adaptively assigned global reward share, driving end-to-end collaboration toward the global objective. Extensive experiments on long-term personalized conversational tasks show CoMAM outperforms leading memory baselines, validating the efficacy of our collaborative reinforcement learning for jointly optimizing multi-agent memory systems.
@@ 2 Related work
Memory System for Personalized LLMs
@@
Optimization of Multi-agent System
Multi-agent Systems (MAS) (Sun et al., 2024; Chen et al., 2024b) leverage multiple specialized agents with heterogeneous configurations and asynchronous execution to tackle complex tasks beyond the capacity of a single agent. To boost MAS performance, existing methods focus on independent optimization, primarily via prompt engineering (Patil et al., 2024; Wu et al., 2024) or fine-tuning of agents on specific tasks (Subramaniam et al., 2025; Yan et al., 2025b). Yet such methods yield only marginal gains, as they overlook the value of joint optimization for inter-agent collaboration. Consequently, recent research has shifted to reinforcement learning (RL)-based collaborative training for end-to-end system optimization, such as MAPoRL (Park et al., 2025) and MARFT (Liao et al., 2025).
Credit Assignment
Credit assignment(Zhou et al., 2020; Bo et al., 2024a; Tseng et al., 2022) concerns properly attributing a team’s global outcome to the specific agents in Multi-Agent Reinforcement Learning (MARL). This is particularly difficult in quantifying agents’ contributions to the outcome, which is often obscured by the actions of other agents and environmental noise. Naive credit assignment approaches evenly distribute global team rewards across agents (Zhao et al., 2026; Hong et al., 2025). Though simple, this leads to significant credit assignment ambiguity: agents cannot distinguish their specific impact, resulting in high variance in policy updates and unstable convergence.
3 Preliminary and Challenges
@@ 3.1 Task Formulation for Multi-agent Memory System
We define the user’s long-term conversational history as $\mathcal{H}=\{h_{1},\dots,h_{N}\}$, which consists of $N$-turn interactions that exceed the LLM’s context window. To facilitate personalization, the Memory System transforms the history $\mathcal{H}$ into multi-granular memories $\mathcal{M}$ that can be retrieved to answer users’ subsequent queries $q$, realized by diverse agents on construction and retrieval tasks (Hu et al., 2025):
  • Memory Construction: The agents (Li et al., 2025) under policies $\{\pi_{\theta}^{\text{cons}_{g}}\}$, distill the raw conversational history $\mathcal{H}$ at different granularities $g$, respectively, finally constructing a multi-granular memory set $M$:
    $\mathcal{M}_{g}=\pi_{\theta}^{\text{cons}_{g}}(\mathcal{H}),$ (1)
    $\mathcal{M}=\{\mathcal{M}_{g}\}.$ (2)
  • Memory Retrieval: Given a subsequent query $q$, retrieval agents (Zhou et al., 2026) under policies $\{\pi_{\theta}^{\text{ret}}\}$ identify the $K$ most relevant memories from $\mathcal{M}$. Then, the agent with policy $\pi_{\theta}^{\text{LLM}}$ generates a response $p$ for the query $q$ using the retrieved $K$ memories. Formally:
    $\{\bm{m}_{k}\}_{k=1}^{K}=\pi_{\theta}^{\text{ret}}(q,\mathcal{M},K),$ (3)
    $p=\pi_{\theta}^{\text{LLM}}\left(q,\{\bm{m}_{k}\}_{k=1}^{K}\right).$ (4)
    For efficiency, $\pi_{\theta}^{\text{ret}}$ and $\pi_{\theta}^{\text{LLM}}$ can be unified into a single policy that performs both memory retrieval and personalized response generation for query $q$.
The global goal of the memory system is to align the generated response $p$ with the user’s personalized preferences and experiences. Formally, given a dataset $\mathcal{D}$ with input pairs $(\mathcal{H},q)$ and a performance evaluation function $\mathcal{F}$, the optimization objective is to find the optimal policy set $\Pi^{*}$ that maximizes the expected global performance:
$\Pi^{*}=\arg\max_{\Pi}\mathbb{E}_{(\mathcal{H},q)\sim\mathcal{D}}\left[\mathcal{F}\left(p^{*},\Pi(\mathcal{H},q)\right)\right].$ (5)
Here, $\Pi=\{\pi_{\theta}^{\text{cons}_{g}},\pi_{\theta}^{\text{ret}}\}$ denotes the policy set, and $p^{*}$ denotes the ground-truth response.
@@ 3.2 Reinforcement Learning for LLMs
To optimize LLMs’ policy $\pi_{\theta}$, Reinforcement Learning (RL) (Ouyang et al., 2022; Zheng et al., 2023) is typically used to maximize the expected reward while maintaining stability relative to a reference model $\pi_{\text{ref}}$. In this work, we primarily build upon Group Relative Policy Optimization (GRPO) (Shao et al., 2024), an efficient RL algorithm that eliminates the need for a separate value function (critic) through group-based relative rewards. Given the input $x$, GRPO samples a group of outputs $\left\{o_{1},\ldots,o_{G}\right\}$ from the current policy $\pi_{\theta}$ and optimizes the objective function:
$\mathcal{J}_{GRPO}\left(\pi_{\theta}\right)=$ (6)
$\frac{1}{G}\sum_{i=1}^{G}\frac{1}{\left|o_{i}\right|}\sum_{t=1}^{\left|o_{i}\right|}\left\{\min\left[\rho_{i,t}\hat{A}_{i,t},\hat{\rho}_{i,t}\hat{A}_{i,t}\right]-\beta\mathbb{D}_{KL}\right\},$
@@
$\hat{A}_{i,t}=\frac{r_{i}-\text{mean}(r)}{\text{std}(r)}.$ (7)
The KL divergence is defined as:
$\mathbb{D}_{KL}=\frac{\pi_{ref}(o_{i,t}|x,o_{i,<t})}{\pi_{\theta}(o_{i,t}|x,o_{i,<t})}-\log\frac{\pi_{ref}(o_{i,t}|x,o_{i,<t})}{\pi_{\theta}(o_{i,t}|x,o_{i,<t})}-1.$ (8)
@@ 4 Method
Figure 3: The overview of our proposed framework, CoMAM, which regularizes agents’ execution as MDP trajectories for joint RL optimization and fosters collaboration via adaptive credit assignment to achieve local-global alignment.
In this section, we present our proposed CoMAM, a Collaborative Reinforcement Learning Framework designed to address the key challenges of jointly optimizing the multi-agent memory systems, as shown in Figure 3. We begin by defining the task-specific agents alongside the local and global rewards in Sec 4.1. Then, we elaborate on regularizing the execution trajectory as a Markov decision process (MDP) in Sec 4.2, which enables joint optimization despite agent heterogeneity and asynchrony. Finally, we present the adaptive credit assignment mechanism in Sec 4.3, which quantifies individual agent contributions and aligns local improvements with global performance.
@@ 4.1 Task-Specific Agents Definition
Building on existing multi-agent memory paradigms (Hu et al., 2025; Li et al., 2025), we define a standard memory system consisting of three specialized agents: Extraction Agent, Profile Agent, and Retrieval Agent.
@@
where $r_{\text{cons}_{f}}$ balances the coverage of relevant information and the filtering of irrelevant content. To prioritize coverage, $\alpha$ is set to $0.8$.
@@
where $r_{\text{cons}_{c}}$ evaluates the quality of $\mathcal{M}_{c}$ relative to the conversation history $\mathcal{H}$ according to predefined rubrics (e.g., scenario, preference, and user profile). $\mathcal{V}$ denotes a frozen pre-trained LLM for reward generation. Finally, the comprehensive memory set is defined as $\mathcal{M}=\{\mathcal{M}_{f},\mathcal{M}_{c}\}$.
Retrieval Agent retrieves top-$k$ relevant memories $\mathcal{E}_{q}^{\prime}=\{\bm{m}_{k}\}_{k=1}^{K}$ from memory set $\mathcal{M}$ and generate personalized response $p$ for user’s subsequent query $q$, guided by policy $\pi_{\theta}^{\text{ret}}$. To evaluate the retrieval performance, we define a rule-based reward function that enables precise matching between the recalled $\mathcal{E}_{q}^{\prime}$ and ground-truth evidence $\mathcal{E}_{q}$:
@@
where $r_{\text{ret}}$ balances the coverage and precision of retrieval. To prioritize retrieval precision while minimizing noise, $\beta$ is set to $0.2$.
To evaluate the memory system’s global performance, we define the rule-based reward for $\pi_{\theta}^{\text{ret}}$ to evaluate its answering accuracy for query $q$. Formally, let $p^{*}$ denote the ground-truth response for the personalized query $q$, we have:
@@
4.2 Execution Trajectory Regularization
Usually, construction agents run periodically to manage historical conversations, while retrieval agents are activated for incoming queries. To enable joint training under such heterogeneous and asynchronous settings, we model the entire construction and retrieval pipeline of the multi-agent memory system as a sequential Markov Decision Process (MDP) $(\mathcal{S},\mathcal{A},T,R)$, where each step is executed by a heterogeneous agent. Formally, let $\pi_{\theta}^{0}=\pi_{\theta}^{\text{cons}_{f}}$, $\pi_{\theta}^{1}=\pi_{\theta}^{\text{cons}_{c}}$$\pi_{\theta}^{2}=\pi_{\theta}^{\text{ret}}$ denote the policies at step $n=1,2,3$, respectively. Then, for each MDP trajectory $\tau$, we have:
  • $\mathcal{S}$: State space, capturing the system state at each step in the MDP. Specifically, $s_{0}=\mathcal{H}$ denotes the Extraction Agent’s initial state of raw conversational history $\mathcal{H}$. $s_{1}=\mathcal{M}_{f}$ denotes the Profile Agent’s state, where $\mathcal{M}_{f}$ is the fine-grained memory extracted by the prior Extraction Agent. $s_{2}=(\mathcal{M},q)$ denotes the Retrieval Agent’s state, with $\mathcal{M}=\{\mathcal{M}_{f},\mathcal{M}_{c}\}$ as the multi-granular memory set and $q$ as the incoming user query. $s_{3}=(\mathcal{E}_{q}^{\prime},p)$ denotes the system’s final state, comprising retrieved information $\mathcal{E}_{q}$ and generated response $p$.
  • $\mathcal{A}$: Action space, defining executable actions of each heterogeneous agent in the MDP. Specifically, $a_{0}$ denotes extracting fine-grained memory by the Extraction Agent. $a_{1}$ denotes abstracting coarse-grained memory by the Profile Agent. $a_{2}$ denotes the Retrieval Agent’s action, which involves retrieving top-$k$ relevant memories and generating a personalized response.
  • $T$: State transition function, where $s_{n+1}=T(s_{n},a_{n})$ and $a_{n}=\pi_{\theta}^{n}(s_{n})$. Specifically, the Extraction Agent transitions to the Profile Agent via $s_{1}=T(s_{0},a_{0})$ and $a_{0}=\pi_{\theta}^{0}(\mathcal{H})$; the Profile Agent then transitions to the Retrieval Agent through $s_{2}=T(s_{1},a_{1})$ and $a_{1}=\pi_{\theta}^{1}(\mathcal{M}_{f})$; finally, the Retrieval Agent produces the system’s final state $s_{3}=T(s_{2},a_{2})$ where $a_{2}=\pi_{\theta}^{2}(\mathcal{M},q)$.
  • $R$: Reward function, quantifying the rewards of agent actions in the MDP. As detailed in Sec 4.1, the local task rewards for each step are $r_{0}=r_{\text{cons}_{f}}$ for the Extraction Agent, $r_{1}=r_{\text{cons}_{c}}$ for the Profile Agent, and $r_{2}=r_{\text{ret}}$ for the Retrieval Agent. And, the global system reward is defined as $r_{3}=r_{\text{ans}}$.
Within this formulation, CoMAM embeds inter-agent dependencies into state transitions along the optimization trajectory. The multi-agent memory system operates sequentially to generate local and global rewards, enabling end-to-end RL optimization with simultaneous updates of heterogeneous policies. The task-specific capabilities and inter-agent dependencies learned in this MDP enable agents to perform well in more complex inference environments, where divergent operating frequencies demand enhanced collaboration.

### 27 Apr 2026 v2

# Joint Optimization of Multi-agent Memory System

@@ Abstract
Memory systems are critical for LLMs, mitigating context window limitations and supporting long-horizon user–LLM interactions. Such systems typically comprise multiple agents responsible for memory construction and retrieval. Existing approaches often optimize each agent independently under a shared global objective (e.g., downstream QA accuracy), treating other agents as a static environment. However, this design has two key limitations: (1) independent optimization ignores inter-agent dependencies and lacks agents’ co-adaptation, and (2) relying solely on sparse global rewards provides limited guidance for optimizing specialized agents and causes ambiguous credit assignment. These may ultimately limit agents’ effective collaboration in the memory system. To address these limitations, we propose CoMAM, a joint optimization framework that promotes collaboration among agents via end-to-end reinforcement learning and an adaptive credit assignment mechanism. Specifically, we model the multi-agent pipeline as a Markov decision process (MDP) to expose inter-agent dependencies during end-to-end training. Agents are then jointly optimized using a combination of their local task reward and an adaptively weighted global reward, enabling agents to co-adapt while receiving targeted feedback for their respective roles. Experiments show that CoMAM consistently outperforms leading memory systems, validating the effectiveness of the joint optimization framework.
@@ 1 Introduction
Memory systems Mem0 ; memorybank ; hu2025memory have emerged as a pivotal component for large language models (LLMs), enabling them to preserve long-horizon conversational history beyond the context window and selectively incorporate it for generation. Such systems are typically implemented as multi-agent pipelines, where construction agents summarize and store user history into multi-granular memories (e.g., long/short or coarse/fine) li2025cam ; mirix ; memory-alpha , and retrieval agents select and integrate relevant memories for downstream queries via RAG RAG or search-based reasoning memagent ; search-r1 . These agents are often instantiated with heterogeneous configurations to handle tasks of varying complexity, enabling flexible modular design and specialization. To enhance the performance of memory systems, existing approaches mem-r1 ; mem1 ; GAM ; memory-alpha often optimize each agent independently via reinforcement learning under a shared global objective (e.g., , QA accuracy on downstream tasks). They typically update one agent at a time, treating the others as fixed components of the environment.
However, such independent optimization has two limitations and may lead to sub-optimal division of labor and collaboration local_optimalty ; Optimas . First, ignorance of dependency among agents: each agent is optimized while treating others as a static environment, preventing the adaptation to other agents’ evolving policies. Second, ambiguous credit assignment: the shared global objective provides insufficient signal to attribute performance improvements to individual agents, which may hinder effective specialization. As illustrated in Figure 1, although agents can be well trained independently (rewards converged in the left figure), combining them directly leads to inferior system performance compared to joint optimization across different context lengths (lower QA accuracy in the middle figure). This gap highlights that independent optimization with the same global system objective is insufficient to achieve optimal system performance, motivating us to jointly optimize these heterogeneous agents in the memory system to facilitate collaboration.
Figure 1: (Left) Training curves of coarse/fine memory construction and retrieval agents, which are optimized independently. (Middle) Global system performance under independent and joint optimization across different context lengths. (Right) Conceptual comparison between existing methods that optimize agents independently with shared global rewards and our joint optimization with adaptive credit assignment.
@@
To address these limitations, we propose CoMAM, a joint optimization framework that promotes collaboration among agents via end-to-end reinforcement learning, together with an adaptive credit assignment mechanism. As illustrated in the right panel of Figure 1, our approach differs from independent optimization by enabling simultaneous updates of heterogeneous agents with adaptively assigned rewards. Specifically, to formulate a standard memory system for coarse/fine memory construction and memory retrieval, we design an extraction agent to construct fine-grained memories (e.g., key events and facts), a profile agent to summarize coarse-grained memories (e.g., users’ stable preferences), and a retrieval agent to retrieve memories and answer queries. To model the inter-agent dependencies for co-adaption, we frame the heterogeneous multi-agent pipeline as a sequential Markov decision process (MDP), enabling agents to coordinate through their intermediate outputs. To alleviate ambiguous credit assignment during joint optimization, CoMAM assigns each agent a combination of its local task reward and an adaptively weighted global reward, providing more targeted feedback for different roles, driving collaboration for the global objective while maintaining task specialization. Extensive experiments on long-horizon conversational tasks show CoMAM outperforms leading memory baselines, validating the efficacy of our joint optimization framework and adaptive credit assignment for multi-agent memory systems.
@@ 2 Related work
Memory Systems for LLMs.
@@
Optimization of Multi-agent Systems.
Multi-agent systems (MAS) masrl_syrvey ; self_play ; Optimas leverage multiple specialized agents with heterogeneous configurations to tackle complex tasks beyond the capacity of a single agent. Existing approaches typically optimize agents independently, either through prompt engineering Gorilla ; avatar or task-specific fine-tuning independent_mas ; mem-r1 . Such independent optimization treats other agents as part of the environment and does not explicitly account for their inner dependence during training. Consequently, recent research has shifted to RL-based joint training for end-to-end system optimization, such as MAPoRL MAPORL and MARFT MARFT .
Credit Assignment.
Credit assignment credit_assign3 ; credit_assign5 ; credit_assign4 concerns properly attributing a team’s global outcome to the specific agents in Multi-Agent Reinforcement Learning. This is challenging due to the interdependence among agents and environmental noise, which obscures each agent’s contribution. A common strategy is to distribute global rewards uniformly across agents naive_assign ; hong2025multi . While simple, this leads to significant credit assignment ambiguity: agents cannot distinguish their specific contributions COMA ; credit_assignment3 , resulting in high variance in policy updates and unstable convergence.
3 Preliminary
@@ 3.1 Task Formulation for Multi-agent Memory System
We define the long-horizon user-LLM conversational history as $\mathcal{H}=\{h_{1},\dots,h_{N}\}$, which consists of $N$-turn interactions that exceed the LLM’s context window. The Memory System transforms the history $\mathcal{H}$ into multi-granular memories $\mathcal{M}$ that can be retrieved to answer users’ subsequent queries $q$, realized by heterogeneous agents on construction and retrieval tasks hu2025memory :
  • Memory Construction: The agents li2025cam under policies $\{\pi_{\theta}^{\text{cons}_{g}}\}$, distill the raw conversational history $\mathcal{H}$ at different granularities $g$, finally constructing a multi-granular memory set $M$:
    $\mathcal{M}_{g}=\pi_{\theta}^{\text{cons}_{g}}(\mathcal{H}),\quad\mathcal{M}=\{\mathcal{M}_{g}\}.$ (1)
  • Memory Retrieval: Given a subsequent query $q$ of downstream task, retrieval agents mem1 under policies $\{\pi_{\theta}^{\text{ret}}\}$ identify the $K$ most relevant memories from $\mathcal{M}$. Then, the agent with policy $\pi_{\theta}^{\text{LLM}}$ generates a response $p$ for the query $q$ using the retrieved $K$ memories. Formally:
    $\{\bm{m}_{k}\}_{k=1}^{K}=\pi_{\theta}^{\text{ret}}(q,\mathcal{M},K),\quad p=\pi_{\theta}^{\text{LLM}}\left(q,\{\bm{m}_{k}\}_{k=1}^{K}\right).$ (2)
    For efficiency, $\pi_{\theta}^{\text{ret}}$ and $\pi_{\theta}^{\text{LLM}}$ can be unified into a single policy that performs both memory retrieval and response generation for query $q$.
The goal of the memory system is to optimize downstream task performance. Given a dataset $\mathcal{D}$ of input pairs $(\mathcal{H},q)$ and an evaluation function $\mathcal{F}$, we aim to learn an optimal policy set $\Pi^{*}$ that maximizes the expected performance:
$\Pi^{*}=\arg\max_{\Pi}\mathbb{E}_{(\mathcal{H},q)\sim\mathcal{D}}\left[\mathcal{F}\big(p^{*},\Pi(\mathcal{H},q)\big)\right].$ (3)
Here, $\Pi=\{\pi_{\theta}^{\text{cons}_{g}},\pi_{\theta}^{\text{ret}}\}$ denotes the policy set, and $p^{*}$ is the ground-truth response.
@@ 3.2 Reinforcement Learning for LLMs
To optimize LLMs’ policy $\pi_{\theta}$, Reinforcement Learning (RL) RLHF ; PPO is typically used to maximize the expected reward while maintaining stability relative to a reference model $\pi_{\text{ref}}$. In this work, we primarily build upon Group Relative Policy Optimization (GRPO) GRPO , an efficient RL algorithm that eliminates the need for a separate value function (critic) through group-based relative rewards. Given the input $x$, GRPO samples a group of outputs $\left\{o_{1},\ldots,o_{G}\right\}$ from the current policy $\pi_{\theta}$ and optimizes the objective function:
$\mathcal{J}_{GRPO}\left(\pi_{\theta}\right)=\frac{1}{G}\sum_{i=1}^{G}\frac{1}{\left|o_{i}\right|}\sum_{t=1}^{\left|o_{i}\right|}\left\{\min\left[\rho_{i,t}\hat{A}_{i,t},\hat{\rho}_{i,t}\hat{A}_{i,t}\right]-\beta\mathbb{D}_{KL}\right\},$ (4)
@@
$\hat{A}_{i,t}=\frac{r_{i}-\text{mean}(r)}{\text{std}(r)},\text{ and }\quad\mathbb{D}_{KL}=\frac{\pi_{ref}(o_{i,t}|x,o_{i,<t})}{\pi_{\theta}(o_{i,t}|x,o_{i,<t})}-\log\frac{\pi_{ref}(o_{i,t}|x,o_{i,<t})}{\pi_{\theta}(o_{i,t}|x,o_{i,<t})}-1.$ (5)
@@ 4 Method
Figure 2: The overview of our joint optimization framework for the multi-agent memory system, CoMAM, which promotes collaboration among agents via end-to-end reinforcement learning and an adaptive credit assignment mechanism.
In this section, a joint optimization framework that enables agents to co-adapt via end-to-end reinforcement learning and an adaptive credit assignment mechanism, as shown in Figure 2. We begin by defining the task-specific agents alongside the local and global rewards in Sec 4.1. Then, we frame the heterogeneous multi-agent pipeline as a Markov decision process (MDP) in Sec 4.2, which models the inter-agent dependencies for joint training. Finally, we present the adaptive credit assignment mechanism in Sec 4.3, which quantifies individual agent contributions and provides more targeted feedback for different agents.
@@ 4.1 Task-Specific Agents Definition
Building on existing multi-agent memory paradigms hu2025memory ; li2025cam , we define a standard memory system consisting of three specialized agents: Extraction Agent, Profile Agent, and Retrieval Agent.
@@
where $r_{\text{cons}_{f}}$ balances the coverage of relevant information and the filtering of irrelevant content. To prioritize fact coverage during the memory construction stage, $\alpha$ is set to $0.8$.
@@
where $r_{\text{cons}_{c}}$ evaluates the quality of $\mathcal{M}_{c}$ relative to the conversation history $\mathcal{H}$ according to predefined rubrics (e.g., scenario, preference, and user profile). More details are presented in Appendix B.3. $\mathcal{V}$ denotes a frozen pre-trained LLM for reward generation. Finally, the comprehensive memory set is defined as $\mathcal{M}=\{\mathcal{M}_{f},\mathcal{M}_{c}\}$.
Retrieval Agent retrieves top-$k$ relevant memories $\mathcal{E}_{q}^{\prime}=\{\bm{m}_{k}\}_{k=1}^{K}$ from memory set $\mathcal{M}$ and generate response $p$ for user’s subsequent query $q$, guided by policy $\pi_{\theta}^{\text{ret}}$. To evaluate the retrieval performance, we define a rule-based reward function that enables precise matching between the recalled $\mathcal{E}_{q}^{\prime}$ and ground-truth evidence $\mathcal{E}_{q}$:
@@
where $r_{\text{ret}}$ balances the coverage and precision of retrieval. To prioritize retrieval precision while minimizing noise for answer generation, $\beta$ is set to $0.2$.
To evaluate the memory system’s global performance, we define the rule-based reward for $\pi_{\theta}^{\text{ret}}$ to evaluate its answering accuracy for query $q$. Formally, let $p^{*}$ denote the ground-truth response for the query $q$, we have:
@@
4.2 Inter-agent Dependency Modeling
In a typical memory system, construction agents first convert conversational histories into structured memories, which are then consumed by retrieval agents to answer queries. This naturally forms a sequential pipeline for heterogeneous agents to coordinate through their intermediate outputs RL_MDP . To capture such inter-agent dependencies during co-adaptation, we model the pipeline as a sequential Markov decision process (MDP), enabling joint optimization across heterogeneous agents. This formulation serves as a training-time abstraction: it does not assume synchronous deployment, but provides a structured way to expose how upstream decisions influence downstream behavior. Formally, the MDP is defined as $(\mathcal{S},\mathcal{A},T,R)$, where each step corresponds to a heterogeneous agent executing its policy $\{\pi_{\theta}^{n}\}_{n=0}^{2}$ (extraction, profile, retrieval). Given an input $(\mathcal{H},q)$, the system evolves as a trajectory $\tau=(s_{0},a_{0},s_{1},a_{1},s_{2},a_{2})$:
  • $\mathcal{S}$: The state represents intermediate outputs of each agent in the MDP, with $s_{0}=\mathcal{H}$, $s_{1}=\mathcal{M}_{f}$, $s_{2}=(\mathcal{M},q)$, and $s_{3}=(\mathcal{E}_{q}^{\prime},p)$.
  • $\mathcal{A}$: Each action $a_{n}$ corresponds to the execution of agent $\pi_{\theta}^{n}$, including fine-grained memory extraction, coarse-grained abstraction, and retrieval with response generation.
  • $T$: The transition function $s_{n+1}=T(s_{n},a_{n})$ propagates intermediate outputs, where each agent’s output becomes part of the next agent’s input, explicitly exposing inter-agent dependencies.
  • $R$: Each step is associated with a local task reward $r_{n}$, together with a shared global reward $r_{3}$ reflecting downstream performance.
Within this formulation, CoMAM embeds inter-agent dependencies into state transitions along the optimization trajectory. The multi-agent memory system operates sequentially to generate local and global rewards, enabling end-to-end RL for heterogeneous policies’ co-adaptation.